OpenStudy (anonymous):

Please help: a gun with a barrel of .63 meters fires a bullet that has a muzzle velocity of 250 m/s. The mass of the bullet is 0.025 kg. What is the force acting on the bullet while it is in the barrel of the gun? If the gun accelerated backward at a rate of 0.75 m/s^2, what is its mass?

OpenStudy (anonymous):

alright lets look at the bullet first, and we are just going to try and figure it out with the units before we do the math. So we know we need to figure out a force so we can apply newton's 3rd law to the gun. We know that force is measured in newtons so lets break down that unit and see what we can come up with. \[N= Kg*\frac{m}{s^2}\]

OpenStudy (anonymous):

So we have a Kg but a lacking a acceleration so I checked out the kinematic equation's and think I found one we have the information to solve for acceleration. \[V_f^2=V_i^2+2ad\]

OpenStudy (anonymous):

Now that we have an acceleration for the bullet we can use the general equation F=ma to calculate the mass of the gun. \[m_{gun}a_{gun}=m_{bul}a_{bul} \rightarrow m_{gun}=\frac{m_{bul}a_{bul}}{a_{gun}}\]

OpenStudy (anonymous):

Then you can check the units

OpenStudy (anonymous):

okay, so can you use Fg=mg for the first step to solve this?

OpenStudy (anonymous):

Or Fnet=ma

OpenStudy (anonymous):

or the acceleration/ velovity equation?

OpenStudy (anonymous):

sorry, it is because I don't think we have learned the kinematic equation.

OpenStudy (anonymous):

We have gone over Newtons 1-3 laws

OpenStudy (anonymous):

I would use the kinematic equation because it asks for the force acting on the bullet inside the barrel.

OpenStudy (anonymous):

ohh wait, my bad.. we have used the kinematic equions, I just wasnt aware it was called that

OpenStudy (anonymous):

the first equation you are talking about is the Kg/(m/s^2)= N .. Right?

OpenStudy (anonymous):

no Kg* (m/s^2)=N

OpenStudy (anonymous):

okay for acceleration I plugged in 250^2=0^+2a(.63)

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

yes that looks right

OpenStudy (anonymous):

okay then I got 99206 then I plugged it into kg(m/s^2) .025(99206)= 2480N

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

@zbay

OpenStudy (anonymous):

you should get a number like 1240n

OpenStudy (anonymous):

divide thatby the acceration of the gun and you get the mass of the gun

OpenStudy (anonymous):

wait, how did you get 1240?

OpenStudy (anonymous):

@zbay

OpenStudy (anonymous):

I used the kinematic equation to find acceleration then multiplied it by the the mass of the bullet.

OpenStudy (anonymous):

which is what I did, but I got 2480

OpenStudy (anonymous):

Did you multiply (.025*99206)?

OpenStudy (anonymous):

@zbay

OpenStudy (anonymous):

ohhh, is it 1240 because tht is the force of one of the forces, not both of the forces added together?

OpenStudy (anonymous):

okay so what do you do after you get 1240?

OpenStudy (anonymous):

the 1240 is both of the forces due to the newton's 3rd law

OpenStudy (anonymous):

okay, so did I multiply the right numbers? (.025)(99206)?

OpenStudy (anonymous):

@zbay

OpenStudy (anonymous):

you are messing up with the calculator somewhere you should end up with .025(49603)

OpenStudy (anonymous):

to be more detailed it would be F=.025(4903)

OpenStudy (anonymous):

ahh, yes you are right, I messed up earlier in the roblem with my calculator, srry.. okay so what do I d after that?

OpenStudy (anonymous):

\[a=\frac{250^2}{.63(2)}\]

OpenStudy (anonymous):

mass=1403/.75

OpenStudy (anonymous):

okay, so tht would be the mass of??

OpenStudy (anonymous):

the gun it seems rather heavy but the math should be right

OpenStudy (anonymous):

okay.. so then what would be the next step?

OpenStudy (anonymous):

thats it you were looking for the mass of the gun

OpenStudy (anonymous):

oh yes.. oky thank you very much! just one more question, if we were to draw a force diagram, what would the force acting on the bullet while it is in the barrel be NAMED?

OpenStudy (anonymous):

@zbay

OpenStudy (anonymous):

do you understand what I am trying to say? do you need me to clarify?

OpenStudy (anonymous):

i'm not sure

OpenStudy (anonymous):

do you know how the force diagram would look like?

OpenStudy (anonymous):

no sorry

OpenStudy (anonymous):

Alright, its ok.. thank you so much for ALL your help! im sorry it took me a while but I really appreciate it, thanks(:

OpenStudy (anonymous):